Melissa invested a sum of money at 3% annual interest. She invested three times that sum at 5% annual interest. If her total yearly interest from both investments was $540, how much was invested at 3%?
Answer by checkley77(12569)

.03X+.05*3X=540
.03X+.15X=540
.18X=540
X=540/.18X
X=3,000 AMOUNT INVESTED @ 3%.
3*3,000=9,000 AMOUNT INVESTED @ 5%.
PROOF:
.03*3,000+.05*9,000=540
90+450=540
540=540